### v1

# A Single-Step Non-Autoregressive Automatic Speech Recognition Architecture with High Accuracy and Inference Speed

Abstract
Non-autoregressive (NAR) automatic speech recognition (ASR) models predict tokens independently and simultaneously, bringing high inference speed. However, there is still a gap in the accuracy of the NAR models compared to the autoregressive (AR) models. To further narrow the gap between the NAR and AR models, we propose a single-step NAR ASR architecture with high accuracy and inference speed, called EfficientASR. It uses an Index Mapping Vector (IMV) based alignment generator to generate alignments during training, and an alignment predictor to learn the alignments for inference. It can be trained end-to-end (E2E) with cross-entropy loss combined with alignment loss. The proposed EfficientASR achieves competitive results on the AISHELL-1 and AISHELL-2 benchmarks compared to the state-of-the-art (SOTA) models. Specifically, it achieves character error rates (CER) of 4.26%/4.62% on the AISHELL-1 dev/test dataset, which outperforms the SOTA AR Conformer with about 30x inference speedup.
keywords:
ASR, E2E, non-autoregressive, single-step NAR, EfficientASR
1 Introduction
Over the past few years, end-to-end (E2E) models have achieved remarkable results on ASR tasks. There are three popular E2E approaches: connectionist temporal classification (CTC) [1], recurrent neural network transducer (RNN-T) [2, 3], and attention-based encoder-decoder (AED) [4, 5]. Among them, the CTC model is non-autoregressive. Although NAR models do not perform as well as AR models due to the conditional independence assumption in decoding, their parallel decoding can greatly reduce the decoding time. To bridge the gap between NAR and AR models, much work has been done.
Depending on the number of iterations in inference, NAR models are classified as iterative or single-step models. Among the former, A-FMLM [6], a non-autoregressive Transformer model (NAT), is the first attempt to introduce the conditional masked language model (CMLM) [7] into ASR. A-FMLM is designed to predict masked tokens conditioned on unmasked ones and whole speech embeddings. However, its target token length needs to be pre-defined, which limits the performance of the model. To address this issue, ST-NAT [8] uses a CTC module to predict the length of the target sequence. Unlike ST-NAT, Mask-CTC and its variants [9, 10, 11] propose to use the CMLM decoder to refine CTC decodings.
However, iterative models require multiple iterations to achieve a competitive result, limiting the speed of inference in practice. To overcome this limitation, single-step NAR models are proposed. LASO [12] implements parallel decoding of Transformer-style models based on the position-dependent summarizer (PDS) module, but it requires a predefined token length. InterCTC [13, 14] uses an additional intermediate loss of CTC to relax the conditional independence assumption of CTC models, improving the accuracy while maintaining the inference speed of the CTC model. CASS-NAT and its variants [15, 16, 17, 18] conduct in-depth exploration on the combination of CTC alignment and NAT, further improving the performance of the single-step NAR models. Different from the CTC-based or NAT-based NAR models mentioned above, Continuous Integrate-and-Fire (CIF) [19] is the first attempt to model ASR tasks by explicitly predicting the length of the target sequence. However, due to the conditional independence assumption, the accuracy of these mentioned NAR models is significantly inferior to the SOTA AR models.
Paraformer [20], a CIF-based model, designs a glancing language model (GLM) based sampler to strengthen the NAR decoder with the ability to model token interdependency. According to the Paraformer’s report, it can achieve comparable performance to that of the SOTA AR transformer on open source Mandarin Chinese datasets [20]. Despite the success of Paraformer in NAR ASR modeling, it requires a well-designed two-pass training process, which makes the model less compact and less efficient.
Inspired by EfficientTTS [21], a NAR Text-to-Speech (TTS) model, we propose a single-step NAR ASR architecture, termed EfficientASR. The EfficientASR is more compact compared to other high-performance NAR models, with efficient training and inference processes. And it achieves fairly competitive results on the public AISHELL-1 and AISHELL-2 benchmarks compared to the SOTA models.
2 Methods
The main contribution of this paper is to introduce the methods from EfficientTTS [21] into NAR ASR modeling. For the ASR task, we redesigned the model structure of EfficientTTS [21] and proposed EfficientASR. In this section, we will illustrate the structure of the proposed EfficientASR.
Figure 1: The model architecture of the proposed EfficientASR. ($\bigotimes$ represents the matrix multiplication operation)
2.1 Overview
The architecture of the proposed EfficientASR is shown in Fig. 1. The training network consists of six modules, namely mel-encoder, text-encoder, alignment predictor, alignment generator, attention reconstruction, and decoder. The text-encoder and alignment generator are designed for training and are no longer activated during inference.
Let the acoustic feature sequence be $\bm{x}\in\mathcal{R}^{T}$ and the transcription sequence be $\bm{y}\in\mathcal{R}^{L}$, where $T$ and $L$ are the length of acoustic feature sequence and transcription sequence respectively. During training, the mel-encoder encodes $\bm{x}$ into the acoustic embeddings $\bm{e}^{(\mathrm{s})}\in\mathcal{R}^{T,d}$, and the text-encoder encodes $\bm{y}$ into text embeddings $\bm{e}^{(\mathrm{t})}\in\mathcal{R}^{L,d}$, where $d$ is the embedding dimension. The alignment generator leverages $\bm{e}^{(\mathrm{s})}$ and $\bm{e}^{(\mathrm{t})}$ to generate the alignment $\bm{\delta}\in\mathcal{R}^{T}$ between acoustic and text embeddings. $\bm{\delta}$ is then passed to the attention reconstruction module to construct an attention matrix $\bm{\hat{\alpha}}\in\mathcal{R}^{L,T}$. Multiplying by $\bm{\hat{\alpha}}$, the acoustic embeddings $\bm{e}^{(\mathrm{s})}$ are transformed into semantic encodings of length equal to the output tokens. The semantic encodings are finally fed into a decoder to predict tokens $\bm{y}^{*}\in\mathcal{R}^{L}$. However, at the inference stage, the model cannot obtain alignments through the generator because of the lack of transcriptions. Therefore, during training, we use the predictor to learn the alignment produced by the generator. Then during inference, the output $\bm{\delta^{*}}$ of the alignment predictor is used to reconstruct the attention matrix and perform the inference of the model.
The mel-encoder is the same as the AR Conformer encoder [22], consisting of several conformer blocks. The text-encoder and the decoder are built with Transformer [23] encoder blocks. Other modules will be detailed in the following parts.
Figure 2: Visualization of alignment generator and attention reconstruction calculation process. (Operation $\Delta$ is defined as Eq. (3). Operation Cum-Sum is defined as Eq. (5).)
2.2 Alignment generator
The alignment generator is built based on the Index Mapping Vector (IMV), proposed in [21], which provides a creative way to describe the alignment between input and output sequences. The calculation process of IMV-based alignment for ASR is depicted in Fig. 2(a). First, a scaled dot-product attention matrix $\bm{\alpha}\in\mathcal{R}^{T,L}$ is built by Eq. (1), with $\bm{e}^{(\mathrm{s})}$ as Query and $\bm{e}^{(\mathrm{t})}$ as Key:
$\alpha_{i,j}=\frac{\mathrm{exp}((\bm{e}_{i}^{(\mathrm{s})}\cdot\bm{e}_{j}^{(\mathrm{t})})*d^{-0.5})}{\sum_{j=0}^{L-1}\mathrm{exp}((\bm{e}_{i}^{(\mathrm{s})}\cdot\bm{e}_{j}^{(\mathrm{t})})*d^{-0.5})}$ (1)
where $10\leq i\leq T-$ 1, $10\leq j\leq L-$ 1, and $d$ is the dimension of $\bm{e}^{(\mathrm{s})}$ and $\bm{e}^{(\mathrm{t})}$. Then the IMV $\bm{p}$ is calculated by Eq. (2):
$\bm{p}=\bm{\alpha}\cdot\bm{I}$ (2)
where $\bm{I}=[0,1,...,L-1]$ is the index vector of transcription, and $\Sigma_{j=0}^{L-1}\alpha_{i,j}=1$. $p_{i}=\Sigma_{j=0}^{L-1}(\alpha_{i,j}*j)$, which is the weighted sum of the attention weight between the $i$-th input frame and each output token and the position index of each output token. $p_{i}$ can be seen as the expected alignment location of the $i$-th input frame in the length range of the output text. We define the increment of alignment locations of adjacent audio frames in the transcription as ${\Delta p_{i}}$, which be calculated by Eq. (3). According to the monotonicity assumption in EfficientTTS [21], when there is a monotonic correspondence between audio and transcription, it can be deduced that $\Delta p_{i}\geq 0$, and the proof process can be found in [21].
$\Delta p_{i}=p_{i}-p_{i-1},1\leq i\leq T-1$ (3)
However, at the beginning of training, the model has not been trained well enough to satisfy $\Delta p_{i}\geq 0$ automatically. To ensure this monotonic relationship, $\Delta p_{i}$ is activated by ReLU, as shown in Eq. (4):
$\delta_{i}=\left\{\begin{array}[]{ll}0,i=0\\ {\rm ReLU}(\Delta p_{i}),1\leq i\leq T-1\end{array}\right.$ (4)
where $\delta_{0}$ is set to 00. $\delta_{i}$ can be understood as the weight of each encoder step involved in the calculation of the decoding label.
Table 1: CER(%) and RTF results of our models on AISHELL-1 and AISHELL-2 tasks and comparison with previous work. (†: RTF is evaluated with batchsize of 8, ‡: Model parameters in the inference phase.)
Model LM AISHELL-1 AISHELL-2 Params RTF ($\downarrow$)
dev test test-ios
AR Transformer [12] w/o 6.1 6.6 7.1 67.5M 0.1900
Conformer [24] w/o - 5.21 5.83 46.25M 0.2100
NAR A-FMLM [6] w/o 6.2 6.7 - - 0.2800
[…]
LASO-big [12] w/o 5.9 6.6 6.7 80.0M 0.0040
TSNAT [25] w/o 5.1 5.6 - 87M 0.0185
Improved CASS-NAT [16] w/o 4.9 5.4 - 38.3M 0.0230
AL-NAT [17] w/o 4.9 5.3 - 71.3M 0.0050
Paraformer [20] w/o 4.6 5.2 6.19 46.2M 0.0168
EfficientASR Base (Ours) w/o 4.30 4.66 6.03 43.6M ‡ 0.0058
EfficientASR Large (Ours) w/o 4.26 4.62 5.76 76.0M ‡ 0.0066
2.3 Attention reconstruction
After obtaining alignment $\bm{\delta}$, our goal is to reconstruct an attention matrix, such that the acoustic embeddings $\bm{e}^{(\mathrm{s})}$ of length $T$ can be transformed into semantic encodings of length $L$ by multiplying with the attention matrix. We follow the similar idea of [21] in reconstructing the attention matrix and the construction process of the attention matrix can be summarized as shown in Fig. 2(b). First, a monotonically increasing alignment position vector $\bm{p\mathrm{{}^{\prime}}}$ is generated by accumulating $\delta_{i}$, as shown in Eq. (5). Then, a scaling strategy as shown in Eq. (6) is applied to ensure that the maximum value of $\bm{p^{\prime}}$ is equal to the length of the output tokens (since the token index starts from 0, in practice, take $L-1$).
$p_{i}{{}^{\prime}}=\sum_{m=0}^{i}\delta_{m},0\leq i\leq T-1$ (5)
$\hat{p}_{i}=\frac{p_{i}{{}^{\prime}}-p_{0}{{}^{\prime}}}{p_{T-1}{{}^{\prime}}-p_{0}{{}^{\prime}}}*(L-1),0\leq i\leq T-1$ (6)
Since both $\bm{\hat{p}}$ and transcription index vector $\bm{I}$ are monotonic sequences ranging from 00 to $L-1$, a distance-aware attention reconstruction is performed by Eq. (7):
$\hat{\alpha}_{i,j}=\frac{{\rm exp}(-\sigma^{-2}*d_{i,j})}{\sum_{i=0}^{T-1}{\rm exp}(-\sigma^{-2}*d_{i,j})}$ (7)
where $d_{i,j}=(\hat{p}_{i}-I_{j})^{2},0\leq i\leq T-1,0\leq j\leq L-1$, represents the semantic distance between the $i$-th encoder frame and the $j$-th token. $\sigma$ is a learnable hyperparameter. Obviously, the smaller distance $d_{i,j}$ is, the larger attention weight $\hat{\alpha}_{i,j}$.
2.4 Alignment predictor
As mentioned above, an alignment predictor is required to generate alignment at inference stage. We built it with two 1D-Convolution layers, each followed by the layer normalization and ReLU function. During training, it takes acoustic embeddings $\bm{e}^{(\mathrm{s})}$ as inputs to generate a predicted alignment $\bm{\delta}^{*}$, while using alignment $\bm{\delta}$ as the label to update the predictor. In the inference stage, we use the alignment $\bm{\delta}^{*}$ produced by the predictor to construct the attention matrix, as shown in the Fig. 1(b). At the same time, we accumulate the $\bm{\delta}^{*}$ and round it to predict the number of output tokens. The Mean Square Error (MSE) loss is used to learn the alignment, and the total loss is defined as:
$\mathcal{L}_{total}=\mathcal{L}_{\rm{CE}}(\bm{y},\bm{y^{*}})+\lambda\mathcal{L}_{\rm{MSE}}(\bm{\delta},\bm{\delta^{*}})\$ (8)
where $\lambda$ is the weight of alignment loss, and $\mathcal{L}_{\rm{CE}}$ is the cross-entropy (CE) loss between the outputs and labels.
3 Experiments
3.1 Experimental Setup
The proposed method is evaluated on two public Mandarin corpus: 178 hours AISHELL-1 [26] and 1000 hours AISHELL-2 [27]. We use 80-channel filter banks computed from a 25 ms window with a stride of 10 ms as features. SpecAugment [28] and speed perturbation are used for data augmentation for all experiments. The sizes of the vocabulary for the AISHELL-1 and AISHELL-2 tasks are $4,233$ and $5,208$, respectively. The EfficientASR model is built in base and large size. For both sizes, the layers of {mel-encoder, text-encoder, alignment predictor, decoder} are set to {12, 1, 2, 6}, while attention head of base model is $4$ with hidden dimension $256$ and attention head of large model is $6$ with hidden dimension $384$. The hyperparameter $\lambda$ in Eq. (8) is set to $1.0$ and the learnable hyperparameter $\sigma$ in Eq. (7) is initialized with $0.5$. The proposed models are developed based on ESPnet [29]  [^1] https://github.com/espnet/espnet.git, and other configurations can be found there. All experiments were performed on $8$ NVIDIA Tesla V100 GPUs.
We use the character error rate (CER) to evaluate the performance of different models and the real-time factor (RTF) to measure the inference speed.
3.2 Results on AISHELL-1 and AISHELL-2
The evaluation results of AISHELL-1 and AISHELL-2 are detailed in Table 1. CER and RTF of previous work are obtained from the paper’s report, except that the results of ESPnet Conformer are obtained from the ESPnet official repository. The RTF of our models is calculated on the AISHELL-1 test set with a batch size of $1$. For a fair comparison with the published work, none of our experiments in Table 1 uses an external LM or pre-training.
For AISHELL-1 task, the proposed EfficientASR models (both base and large size) outperform all other models presented. On the AISEHLL-1 test set, our large model achieves a CER of 4.62%, which is 0.58% (absolutely) better than the previous best result of NAR models (from Paraformer [24]). Furthermore, the inference speed of our large model is twice more than that of Paraformer.
When compared with the AR models, our model not only outperforms SOTA AR Conformer in CER, but also has inference speed 30x faster than AR Conformer, which is an impressive performance. Since AISHELL-1 contains less than 200 hours traning data, we also validate our method on a larger corpus, 1000 hours AISHELL-2. On the AISHELL-2 test ios set, our large model achieves a CER of 5.76%, which also outperforms all AR and NAR models presented.
3.3 Analysis
To more fairly evaluate the performance of our model, we use ESPnet[29] to reproduce the SOTA AR and NAR models, namely Conformer and Paraformer. The results are shown in Table 2. With a similar model size, the base version of EfficientASR achieves the lowest CER and RTF in the Aishell-1 test set, which verifies the effectiveness of the proposed method.
Table 2: Reproduction CER (%) and RTF results of the SOTA AR model Conformer and NAR model Paraformer on the AISHELL-1 task and comparison with our model.
Model AR/NAR dev test RTF ($\downarrow$)
Conformer AR 4.30 4.70 0.3980
Paraformer NAR 4.73 5.11 0.0136
EfficientASR Base NAR 4.30 4.66 0.0058
Table 3: Reproduction Inference time(s) in different modules of Paraformer and EfficientASR Base. Each item is the cumulative inference time on the entire AISHELL-1 test set.
Model Encoder Predictor Decoder
Paraformer 208.76 191.79 56.58
EfficientASR Base 209.55 9.65 32.60
In Table 3, we give a detailed analysis of inference time in different modules of Paraformer and the base version of EfficientASR. Note that here Paraformer’s predictor includes the CIF module, while ours includes the attention reconstruction module. As can been seen, the EfficientASR Base is much faster in the predictor module. This is mainly caused by the CIF mechanism in Paraformer. When decoding, CIF needs to recursively obtain the acoustic boundary of each emitted token in advance, while our model performs parallel calculations during the entire inference process. In addition, our decoder is also faster. This is because Paraformer’s decoder is built using transformer decoder blocks, while our decoder is built using transformer encoder blocks that do not include cross-attention layers.
Table 4: Analysis of decoding results (%). Baseline: the alignment for attention reconstruction is obtained from the predictor. Oracle: the alignment for attention reconstruction is obtained from generator based on ground-truth labels. Note: Using only the Baseline decoding method, we have achieved SOTA CER performance on corresponding benchmarks.
AISHELL-1 AISHELL-2
dev test dev-ios test-ios
Baseline (SER) 2.12 2.43 3.09 3.29
Baseline (CER) 4.26 4.62 5.64 5.76
   + 4-gram 4.06 4.37 5.31 5.40
Oracle (CER) 3.61 3.93 3.78 4.10
In Table 4, we further explore the potential of the proposed model. Chinese characters are monosyllabic characters in which different characters may have the same Pinyin syllable (the phonetic symbols of Mandarin). When we analyze the decoding results, we find that many misrecognized characters have the same syllable as their reference characters, in other words, although these characters are misrecognized, their syllables are correctly recognized. So we use a pronunciation lexicon to map the character-based decoding results to syllable-based decoding results to test the syllable error rate (SER) (e.g. Chinese characters ”{CJK*}UTF8gbsn你 好 ”, and their syllables are ”$ni3\ hao3$”). As is shown in Table 4, SER is significantly lower than CER. This means that if we can eliminate this kind of errors, there is still a lot of room for improvement in CER performance. Therefore, we consider using beam search decoding (the beam size is $10$) combined with a 4-gram LM to reduce such errors. The 4-gram LMs are trained using the training set text of the corresponding corpus. It can be seen that the CER results using 4-gram have been improved to a certain extent. However, there is still a large gap between CER and SER, which means that the proposed EfficientASR still has great potential for improvement in semantic modeling. This provides guidance for our future work.
On the other hand, to evaluate the modeling ability of the alignment predictor, we analyze the gap between Baseline CER and Oracle CER, where Baseline’s alignment is obtained from the alignment predictor and the Oracle’s alignment is obtained from the alignment generator based on the ground-truth labels. As shown in Table 4, the Oracle CER is much lower than the Baseline CER. Since the alignment predictor in our model is decoupled from the generator, the performance of the model can be improved to a new level if the predictor can better learn the alignment produced by the generator, which provides another direction for the optimization of the model.
Figure 3: Alignment plots of both the baseline model and proposed model. The first plot is the alignment plot of the Conformer, while the second plot is the alignment constructed by the proposed EfficientASR. The horizontal axis represents the input frame step, and the vertical axis represents the output step.
To further analyze the behaviour of the proposed model, we plot the alignment matrices comparison between the Conformer and the proposed EfficientASR in Figure 3. The alignment matrix of Conformer is from the last cross-attention layer of the model. As can be seen, the proposed method generates more monotonic and cleaner alignments compared to Conformer.
4 Conclusion
In this paper, we propose a novel single-step NAR ASR architecture called EfficientASR. By leveraging our proposed methods for generating monotonic alignments and constructing alignment matrices, EfficientASR not only possesses a simpler structure but also achieves leading performance in terms of inference efficiency and accuracy. Experiments on the open source dataset AISHELL-1 demonstrate that the proposed EfficientASR achieves comparable CER performance to SOTA AR Conformer, with a 30x decoding speedup. On the other hand, we find that the CER performance of our model has great potential for improvement. Our next work will focus on improving the semantic modeling ability of the proposed EfficientASR. Besides, we will study the performance of the EfficientASR in alphabetic languages (e.g. English).